OpenStudy (anonymous):

Assume x is greater than 0. Simplify sin(2arccos(.5x)). The answer should have no trig functions in it, but it will have an x.

OpenStudy (anonymous):

is that 0.5x?

OpenStudy (anonymous):

yes! sorry for the confusion

OpenStudy (anonymous):

what is is the equivalent form of sin(2u) ?

OpenStudy (anonymous):

2cosxsinx ?

OpenStudy (anonymous):

correct. If you apply the identity to the original expression, what do you get?

OpenStudy (anonymous):

im not sure how to do that

OpenStudy (anonymous):

you get 2 sin(arccos(x/2)) cos(arccos(x/2)), yes?

OpenStudy (anonymous):

Ah ok! got that! Can I then cancel out the cos(arccos(0.5))?

OpenStudy (anonymous):

no, you can not

OpenStudy (anonymous):

you can not do so without stating the domian first

OpenStudy (anonymous):

Oh I just saw the mistake in that. That was dumb. Should I use half angle identities then?

OpenStudy (anonymous):

no you don't have too. What is the domain?

OpenStudy (anonymous):

The domain of arccos and arcsin is -1 to 1

OpenStudy (anonymous):

ok, if the domain of arccos(x) is [-1,1], what is the domain of arccos(x/2) ?

OpenStudy (anonymous):

[-2, 2] ?

OpenStudy (anonymous):

correct. Now, you're allowed to do cos(arccos(x/2)) = x/2. Now you see why you need to state the domain first. If you didn't, cos(arccos(x/2)) = x/2 is false if x is not in [-2,2]

OpenStudy (anonymous):

say x = 10. cos(arccos(10/2)) is undefined, but 10/2 = 5. Undefined is not the same as 5 (which is defined)

OpenStudy (anonymous):

now what about sin(arccos(x/2)) ?

OpenStudy (anonymous):

well i can't cancel them out... Im not sure

OpenStudy (anonymous):

maybe the half-angle identities?

OpenStudy (anonymous):

no, half angle identity would probably make it worse. Let's let u = arccos(x/2), then cos(u) = x/2, yes?

OpenStudy (anonymous):



OpenStudy (anonymous):

so am i solving for u?

OpenStudy (anonymous):

sin(u) = sqrt(4-x^2)/2 but since u = arccos(x/2), sin(u) = sin(arccos(x/2)) = sqrt(4-x^2)/2

OpenStudy (anonymous):

so you have, 2 (sqrt(4-x^2)/2) (x/2) = (x/2) sqrt(4-x^2) <-- ans

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

no prob